OpenStudy (anonymous):

clove the equation: 2/3 (x − 4) = 2x for x.

OpenStudy (anonymous):

solve*

OpenStudy (anonymous):

x-4 = 2x *3/2 x-4 =3x x-3x =4 -2x =4 x=???

OpenStudy (anonymous):

Multiply each term on the left by 2/3 and then divide that answer by 2 so solve for x.